OpenStudy (anonymous):

An soccer ball is launched horizontally off a cliff that is 12 meters tall with in initial velocity of 5 meters per second. The ball experiences a resistant force where the ball decelerates by ax=3 m/s2 in the x direction and ay=3.8 m/s2 in the y direction. How long does the object fall for?

OpenStudy (ivancsc1996):

Well, the deceleration in the x direction doesn't matter for the duration of the fall. If I understand the problem there is an extra deceleration in y yielding the following equation\[0=h-\frac{ 1 }{ 2 }(g-a)t^2 \rightarrow t=\sqrt{\frac{ 2h }{ g-a }}=\sqrt{\frac{2 \times 5m}{9.8\frac{ m }{ s^2 }-3.8\frac{ m }{ s^2 }}}=\sqrt{2.5}s\]

OpenStudy (surry99):

Read the question carefully...h is not 5m.

OpenStudy (ivancsc1996):

Oh sorry exchange 5m for 12m

OpenStudy (ivancsc1996):

t=2s